# Spectra and transport of probability and energy densities in a $\mathcal{PT}$-symmetric square well with a delta-function potential

### v1

@@ III.1.2 Scattering states
As in the case for bound states, we have efficient transport also in this case. Relation (56) is also interesting because from it one can derive a kind of generalized unitary condition. In order to do so, we start with condition (56) and we evaluate Eq. (52) in $x=-b$, this yields
@@
In the case of Hermitian systems, this relation reduces to the standard probability conservation $|\mathbf{t}_{\pm}(k)|^{2}+|\mathbf{r}_{\pm}(k)|^{2}=1$, in our case probability is not conserved in general.
@@ IV Summary and conclusions
Regarding the transport properties of the system, the transmission and reflection coefficients were calculated. As expected, the zeros of the left reflection coefficient correspond to the eigenvalues of the bound states, because the incident wave and the transmitted wave both vanish asymptotically, so at the zeros of the transmission, the scattered and bound wave functions are the same. In the inverse process, the scattering from right to left, the states that correspond to the zeros of the reflection coefficient are a discrete set of reflectionless states that diverge to infinity.
The transport efficiency of the system can be quantified by evaluating the density fluxes at the edges of the well, that is, in the positions of the source and drain. In these regions, we can observe that, everything that comes from the source reaches the sink, the gain and loss balance, such that there is no accumulation or depletion of probability or energy in the system. This characteristic has also been reported for tight-binding Hamiltonians with gain and loss sites Ortega et al. (2020, 2021). However, as the strength of the delta increases it acts as a kind of barrier, which modulates the flow of probability and energy, interfering with the transport efficiency of the system, and leading to a break in the $\mathcal{PT}$-symmetry.
Since the system studied here can be solved exactly, it would be of interest to use our findings in extending the analysis to some other effects which might also be of potential interest. These comprise, but are not limited to, the case of analyzing time-dependent wavepackets (e.g. Gaussian wavepackets) and its evolution throughout the $\mathcal{PT}$ scattering system, the so-called topological energy transfer with exceptional points Xu et al. (2016), the study of the jamming anomaly Barashenkov et al. (2016), etc.
@@ Appendix A Transfer matrix
Is well know that the transfer matrix (and its inverse) can be written in terms of the transmission and reflection coefficients (denoted $\mathbf{r},\mathbf{t}$ respectively), and considering the above we have Mostafazadeh (2018)
@@
which means no reciprocity in transmission. Hence, we write the wave function of scattering process as

### v2

@@ III.1.2 Scattering states
As in the case for bound states, we have efficient transport, in the sense that the total gain and loss in the system balance exactly, also in this case. Relation (56) is also interesting because from it one can derive the generalized unitary condition Eq.(34). In order to do so, we start with condition (56) and we evaluate Eq. (52) in $x=-b$, this yields
@@
In the case of Hermitian systems, this relation reduces to the standard probability conservation $|\mathbf{t}_{\pm}(k)|^{2}+|\mathbf{r}_{\pm}(k)|^{2}=1$.
@@ IV Summary and conclusions
Regarding the transport properties of the system, the transmission and reflection coefficients were calculated. As expected, the zeros of the left reflection coefficient correspond to the eigenvalues of the bound states, because the incident wave and the transmitted wave both vanish asymptotically, so at the zeros of the transmission, the scattered and bound wave functions are the same. In the reverse process, the scattering from right to left, the states that correspond to the zeros of the reflection coefficient are a discrete set of reflectionless states that diverge to infinity.
The transport efficiency of the system can be quantified by evaluating the density fluxes at the edges of the well, that is, in the positions of the source and drain. In these regions, we can observe that, everything that comes from the source reaches the sink, the gain and loss balance, such that there is no accumulation or depletion of probability or energy in the system. This characteristic has also been reported for tight-binding Hamiltonians with gain and loss sites Ortega et al. (2020, 2021). However, as the strength of the delta increases it acts as a kind of barrier, which modulates the flow of probability and energy, interfering with the transport efficiency of the system, and eventually leading to a break in the $\mathcal{PT}$-symmetry. In this sense, unlike other systems, the parameter that modulates the source ($V_{I}$) does not cause a drop in the flow that precedes the appearance of the exceptional point, as has been reported in other systems Barashenkov et al. (2016).
Regarding the analysis of the energy fluxes, we mention that this is a topic addressed at least since the 60s Epstein et al. (1965), and still under discussion in different instances (see Mathews (1974); Wu and Dvira (2009); Arvizu et al. (2023); Stepanyan and Allahverdyan (2024)). Up to now, there is no indication that the matter has been settled, in the sense that a standard definition for the energy density is still missing. We address this quantity here because we believe that experiments based on $\mathcal{PT}$-symmetric systems can be carried out in order to determine the nature of this quantity, and decide how an energy density should be defined in Quantum Physics.
Finally, since the system studied here can be solved exactly, it would be of interest to use our findings in extending the analysis to some other effects which might also be of potential interest. These comprise, but are not limited to, the case of analyzing time-dependent wavepackets (e.g. Gaussian wavepackets) and their evolution throughout the $\mathcal{PT}$ scattering system, the so-called topological energy transfer with exceptional points Xu et al. (2016), the study of the jamming anomaly Barashenkov et al. (2016), etc.
@@ Appendix A Transfer matrix
Is well known that the transfer matrix (and its inverse) can be written in terms of the transmission and reflection coefficients (denoted $\mathbf{r},\mathbf{t}$ respectively), and considering the above we have Mostafazadeh (2018)
@@
which means that there is no reciprocity in transmission from one or the other direction. Hence, we write the wave function of the scattering process as